# SyntheT2C: Generating Synthetic Data for Fine-Tuning Large Language Models on the Text2Cypher Task

### v1

@@ Abstract
Integrating Large Language Models (LLMs) with existing Knowledge Graph (KG) databases presents a promising avenue for enhancing LLMs’ efficacy and mitigating their “hallucinations”. Given that most KGs reside in graph databases accessible solely through specialized query languages (e.g., Cypher), there exists a critical need to bridge the divide between LLMs and KG databases by automating the translation of natural language into Cypher queries (commonly termed the “Text2Cypher” task). Prior efforts tried to bolster LLMs’ proficiency in Cypher generation through Supervised Fine-Tuning. However, these explorations are hindered by the lack of annotated datasets of Query-Cypher pairs, resulting from the labor-intensive and domain-specific nature of annotating such datasets. In this study, we propose SyntheT2C, a methodology for constructing a synthetic Query-Cypher pair dataset, comprising two distinct pipelines: (1) LLM-based prompting and (2) template-filling. SyntheT2C facilitates the generation of extensive Query-Cypher pairs with values sampled from an underlying Neo4j graph database. Subsequently, SyntheT2C is applied to two medical databases, culminating in the creation of a synthetic dataset, MedT2C. Comprehensive experiments demonstrate that the MedT2C dataset effectively enhances the performance of backbone LLMs on the Text2Cypher task. Both the SyntheT2C codebase and the MedT2C dataset will be released soon.
@@ 1 Introduction
Knowledge Graphs (KGs) constitute vital reservoirs of information within the Retrieval-Augmented Generation (RAG) paradigm Lewis et al. (2020) of Large Language Models (LLMs). Distinguished from other information sources, KGs boast structured and meticulously curated data, rendering them conducive to seamless updates and rectifications. Such attributes position KGs as pivotal instruments for mitigating issues of knowledge cutoff and “hallucinations” within LLMs. Notably, KGs have long served as a core in numerous knowledge-intensive products and applications Kertkeidkachorn et al. (2023); Cui et al. (2024); Xu et al. (2020). With the advent of LLMs, many researchers have focused on synergizing KGs with LLMs following the RAG framework. The inherent fidelity and adaptability of KGs make them practical assets for deployment in production environments, and also catapult KGs to the forefront of academic research.
@@
While KGs represent invaluable repositories of reference information, their efficient utilization remains a formidable challenge. Early methodologies involved direct extraction of triplets from KGs, subsequently integrating these text-form triplets directly into the prompts of LLMs Fatemi et al. (2023). However, this approach often fails to concurrently preserve both semantic and structural nuances inherent within the KG. An alternative approach involves querying existing graph databases just like human users, promising accurate and interpretable results. Nonetheless, the primary impediment lies in the LLM’s ability to formulate correct and executable queries. To address this limitation, numerous query generation tools or methodologies Zhang et al. (2022); Abdelaziz et al. (2021); Shen et al. (2023) are proposed, aiming to translate human users’ natural language queries into query languages. This task assumes paramount importance for LLM development for two pivotal reasons: (1) it empowers LLMs to consistently produce reliable queries, thereby augmenting their utilization of existing KG databases to address knowledge deficits; (2) it facilitates human interaction with KG databases through natural language, substantially lowering the barrier to entry for KG database utilization. Among the spectrum of query generation research, the sub-task of translating natural language into the Cypher Francis et al. (2018) query language for Neo4j Neo4j (2012) databases stands out as a prominent research focus. This prominence is attributed to two key factors. Firstly, Neo4j is a widely adopted solution for constructing KG databases, positioning Cypher as an essential tool for accessing these extensive repositories. Secondly, Cypher is a query language specifically designed for querying graph structures, offering significantly faster performance than other query languages, such as SQL, when processing graph data. Consequently, our work centers on this sub-task, commonly termed as “Text2Cypher” (T2C).
@@
A similar task to the Text2Cypher task is the “Text2SQL” task, wherein researchers endeavor to translate natural language sentences into SQL queries. Leveraging manually annotated datasets like SPIDER Yu et al. (2019), numerous methodologies have emerged, including SpCQL Guo et al. (2022) and SQLNet Xu et al. (2017). Conversely, scant attention has been directed towards the Text2Cypher task. Existing approaches typically resort to decomposing the original query into smaller components and translating each part separately. For instance, $R^{3}$-NL2SQL Zhou et al. (2023b) partitions the query generation process into CRUD keywords prediction, clause selection, and object type identification. Despite the success of these methods, adapting them to a specific KG database demands substantial extra effort. With the rise of LLMs, using LLMs for Cypher query generation appears promising. Notably, to the best of our knowledge, no endeavors have explored the potential application of LLMs to the Text2Cypher task. Our work aims to bridge this gap in the literature.
@@
The Cypher writing performance of vanilla LLMs is not satisfactory. To improve it, we employ SFT, which necessitates a dataset of Question-Query pairs. However, creating such a dataset is challenging as it requires both domain-specific knowledge of the KG’s content and expertise in Cypher’s syntax. Consequently, there is currently no annotated dataset for the Text2Cypher task. To overcome this obstacle, we introduce SyntheT2C, a method designed to produce high-quality synthetic Question-Cypher pairs through two distinct pipelines: LLM-based prompting and template-filling (as shown in Figure 1). The LLM-based prompting pipeline aims to generate Cypher queries with greater semantic flexibility, while the template-filling pipeline focuses on producing syntactically complex Cypher queries. The generated Question-Query pairs undergo rigorous automated and manual validation, before being used to fine-tune backbone LLMs. The performance of Cypher generation is evaluated with a manually annotated evaluation dataset, complemented by a qualitative assessment using GPT as a judge. Additionally, we conduct a scalability test by fine-tuning the LLMs with larger synthetic datasets, which demonstrates that the synthetic data generated using our method does not collapse into simple patterns, thereby establishing the robustness of our approach for larger-scale applications.
@@
SyntheT2C is tested with two medical databases: the LHY database and the Hetionet database (details in Section 5). The generated synthetic dataset, “MedT2C”, will be made public.
@@ 2.1 Knowledge Graph and graph database
In recent years, KGs have emerged as fundamental resources for organizing, representing, and querying vast amounts of interconnected information or domain-specific knowledge. These graphs find applications across various domains, including but not limited to, healthcare Cui et al. (2024); Abu-Salih et al. (2022), finance Elhammadi et al. (2020); Kertkeidkachorn et al. (2023), and e-commerce Xu et al. (2020). In the realm of Natural Language Processing and Artificial Intelligence (AI), KGs serve as invaluable sources of context and factual knowledge, enabling systems to reason, infer, and generate responses with enhanced accuracy and coherence.
To handle the processing of graph-structured data, a series of graph databases were invented, including Neo4j Neo4j (2012), NebulaGraph Wu et al. (2022), and Amazon Neptune Bebee et al. (2018). Among them, our work focuses on the Neo4j database Neo4j (2012), a widely used graph database management system that excels in modeling and querying highly interconnected data. Neo4j database employs a powerful query language called Cypher for expressing complex graph patterns and retrieving specific data subsets.
@@ 2.2 Large Language Models
LLMs are advanced AI models that have been trained on vast amounts of text data to understand and generate human-like language. After the recent breakthrough marked by the release of InstructGPT Ouyang et al. (2022) by OpenAI, a series of LLMs are released, featuring different advantages and drawbacks, e.g., the series of GPT models Brown et al. (2020); OpenAI (2023) by OpenAI, Llama Meta (2024) by Meta, Qwen Bai et al. (2023) by Alibaba Cloud, InternLM Cai et al. (2024b) by Shanghai AI Lab, etc. LLMs can comprehend and generate text across a wide range of topics and writing styles. Recent researches highlight their ability to utilize external existing tools like calculator, search engine, or databases Patil et al. (2023); Nakano et al. (2022); Cai et al. (2024a); Qin et al. (2023). This ability is usually abstracted as “Function calling”, and many of its implementations involve generating codes or queries with LLMs to interact with external tools.
@@ 2.3 Code generation
Code Generation is the process of automatically producing executable code from a higher-level representation or natural language. With the advent of LLMs, code generation has experienced a significant advancement. LLMs can now be trained on vast amounts of code and programming-related text materials, enabling them to understand and generate code snippets based on given requirements (e.g., Codex Chen et al. (2021), Polycoder Xu et al. (2022), and Code Llama Rozière et al. (2024)). By leveraging the contextual understanding Dong et al. (2023) and language capabilities of LLMs, code generation becomes more efficient, accurate, and adaptable. Code generation with LLM is not only useful in helping developers to write codes but also in providing a powerful “language” for LLM to interact with other tools: LLMs can be tuned to output executable codes or queries to manipulate external resources. This is the fundamental idea for research in “Function Calling” and Multi-Agent Systems. Current code generation methods rely on two methods for evaluation: either with automatic metrics calculated with an annotated evaluation dataset Papineni et al. (2002); Lin (2004); Banerjee and Lavie (2005); Evtikhiev et al. (2023); Zhou et al. (2023a) or with comparison by a judge (human or powerful LLM like GPT-4) Zheng et al. (2023). Both evaluation methods are used in our work.
@@ 3.2.1 LLM-based prompting pipeline
This pipeline adopts an idea similar to Knowledge Distillation: to SFT a weaker LLM, we could use the Cyphers generated by stronger LLMs. While off-the-shelf LLMs are typically not optimized for Cypher query generation, as Cypher likely constitutes only a small fraction of their pre-training data, they have nonetheless demonstrated strong in-context learning capabilities Dong et al. (2023). Therefore, half of $\mathcal{S}$ is built by few-shot prompting GPT-4o OpenAI (2023). To further simplify the task of generation and to ensure a higher quality of the generated data, we split the whole generation task into (1) extracting information from the database; (2) determining the question categories; and (3) generating the Cyphers for each category with extracted information.
@@ 3.2.2 Template-filling pipeline
One such template is illustrated in Figure 4. In this example, the subschema is introduced to manage cases where the entire database cannot be loaded at once, necessitating the selection and injection of only the relevant subgraph into the prompt. The variables label_i and prop_j represent the randomly sampled names of nodes and their attributes. These templates are initially crafted taking inspiration from Cypher Generator Onofrei (2024), then enriched and verified by the authors.
@@ 3.3 Quality validation
To ensure the quality of the generated synthetic Question-Cypher pairs before their application in SFT, it’s imperative to conduct thorough validation to prevent the “garbage in, garbage out” scenario. However, manually scrutinizing thousands of Cypher queries is arduous and time-consuming. In response, a suite of automatic validators has been implemented to alleviate the burden of manual inspection. In the end, the Cyphers that pass through these automated validators undergo a final round of meticulous manual validation by researchers.
@@ 3.3.1 Automatic validation
The design of Semantic Validator is inspired by the research in machine translation Hoang et al. (2018). This validator utilizes an LLM to translate the generated Cypher back into a natural language question. It then computes the semantic similarity between the translated question and the original question. If the similarity score exceeds a predefined threshold, the Cypher passes validation. We also implement an alternative version of the Semantic Validator, where the LLM assesses semantic similarity directly. Both versions produce coherent validation results, with the latter being adopted for efficiency in subsequent experiments. The prompt used in this validator is presented in Appendix C.1.
The Entity Validator assesses the coverage of entities in the generated Cyphers. The entities in the original question $q$ are extracted via Named Entity Recognition (NER) using the spaCy Honnibal and Montani (2017) model en_core_web_sm. Entities in the generated Cypher $c$ are parsed and extracted using Regular Expressions. A successful validation requires 100% coverage of $q$’s entities in $c$. English entities are first transformed into lemmas using spaCy for fuzzy matching.
@@
Lastly, the Coherence Validator executes the Cypher against $\mathcal{B}$ and evaluates the coherence between the execution results and the original question with LLM, using the prompt presented in Appendix C.2.
@@ 3.3.2 Manual validation
Each Cypher checked by the validators is randomly assigned to two researchers, who independently assess its quality. If both researchers provide a unanimous judgment, their consensus is adopted. In cases of divergent opinions, a third researcher is brought in for further review. The final validation outcome for such Cyphers is determined through a majority vote among the three researchers.
@@ 4.1 LHY and Hetionet Graph databases
Throughout our experiment, we employed two Neo4j databases housing general medical knowledge in graph form: the LHY Medical Knowledge Database (referred to as “LHY”) and the Hetionet Medical Knowledge Database (referred to as “Hetionet”). Both databases are publicly accessible, differing primarily in language: the data within the LHY database is presented in Chinese, whereas Hetionet is written in English.
The LHY Database Liu (2018) serves as the backend database for a Medical Question-Answering system. This database comprises comprehensive medical knowledge, encompassing a wide array of diseases, symptoms, drugs, and related information. Its content is sourced from medical websites, meticulously cleaned, reorganized, and stored within a Neo4j database. There are about 44k entities and 300k relations in it.
Hetionet Himmelstein et al. (2017) is an open and free-to-use database of biomedical knowledge resource implementing “hetnet” model. Aggregating insights from 29 public databases, Hetionet boasts a knowledge network spanning various fields, encompassing a wide array of entities, including genes, compounds, anatomical structures, diseases, symptoms, side effects, etc. There are approximately 47k entities and 2.2 million relations in the Hetionet database.
The detailed statistics of both databases are presented in Appendix D.

### v2

@@ Abstract
Integrating Large Language Models (LLMs) with existing Knowledge Graph (KG) databases presents a promising avenue for enhancing LLMs’ efficacy and mitigating their “hallucinations”. Given that most KGs reside in graph databases accessible solely through specialized query languages (e.g., Cypher), it is critical to connect LLMs with KG databases by automating the translation of natural language into Cypher queries (termed as “Text2Cypher” task). Prior efforts tried to bolster LLMs’ proficiency in Cypher generation through Supervised Fine-Tuning (SFT). However, these explorations are hindered by the lack of annotated datasets of Query-Cypher pairs, resulting from the labor-intensive and domain-specific nature of such annotation. In this study, we propose SyntheT2C, a methodology for constructing a synthetic Query-Cypher pair dataset, comprising two distinct pipelines: (1) LLM-based prompting and (2) template-filling. SyntheT2C is applied to two medical KG databases, culminating in the creation of a synthetic dataset, MedT2C. Comprehensive experiments demonstrate that the MedT2C dataset effectively enhances the performance of backbone LLMs on Text2Cypher task via SFT. Both the SyntheT2C codebase and the MedT2C dataset are released in https://github.com/ZGChung/SyntheT2C.
@@
Zijie Zhong Shanghai AI Laboratory zhongzijie@pjlab.org.cn                        Linqing Zhong Beihang University lqzhong@buaa.edu.cn                        Zhaoze Sun Beihang University szz20241050@buaa.edu.cn
Qingyun Jin Beihang University sy2303108@buaa.edu.cn                        Zengchang Qin [^1] Corresponding authors Beihang University zcqin@buaa.edu.cn                        Xiaofan Zhang [^1] Corresponding authors Shanghai AI Laboratory zhangxiaofan@pjlab.org.cn
@@ 1 Introduction
Knowledge Graphs (KGs) constitute vital reservoirs of information within the Retrieval-Augmented Generation (RAG) paradigm (Lewis et al., 2020) of Large Language Models (LLMs). Distinguished from other information sources, KGs boast structured and meticulously curated data, rendering them conducive to seamless updates and rectifications. Such attributes position KGs as pivotal instruments for mitigating issues of knowledge cutoff and “hallucinations” within LLMs. The inherent fidelity and adaptability of KGs make them practical assets for numerous knowledge-intensive products and applications (Kertkeidkachorn et al., 2023; Cui et al., 2024; Xu et al., 2020). With the advent of LLMs, many researchers have focused on synergizing KGs with LLMs following the RAG framework, catapulting KGs to the forefront of academic research.
@@
The efficient utilization of KG remains a formidable challenge because of the difference in format. Early methodologies involve direct extraction of triplets from KGs, subsequently integrating these text-form triplets directly into the prompts of LLMs (Fatemi et al., 2023). However, this approach often fails to concurrently preserve both semantic and structural nuances inherent within the KG. An alternative approach involves querying existing graph databases just like human users, promising accurate and interpretable results. Nonetheless, the primary impediment lies in the LLM’s ability to formulate correct and precise queries. To address this limitation, numerous query generation tools or methodologies (Zhang et al., 2022; Abdelaziz et al., 2021; Shen et al., 2023) are proposed, aiming to translate human users’ requests (in natural language) into query languages. This task assumes paramount importance for two pivotal reasons: (1) it empowers LLMs to consistently produce reliable queries, thereby enabling them to address knowledge deficits via direct interaction with KG databases; (2) it significantly facilitates human users’ interaction with KG databases because learning the specific query language is no longer necessary. Among the spectrum of query generation research, the sub-task of translating natural language into the Cypher (Francis et al., 2018) query language for Neo4j (Neo4j, 2012) databases stands out as a prominent research focus for two factors: (1) Neo4j is a widely adopted solution for KG databases, positioning Cypher as an essential tool for accessing these extensive repositories; (2) Cypher is a query language specifically designed for querying graph structures, offering significantly faster performance than other query languages, such as SQL, when processing graph data. Consequently, our work centers on this sub-task, commonly termed as “Text2Cypher” (T2C).
@@
A similar task to the Text2Cypher task is the “Text2SQL” task, wherein researchers endeavor to translate natural language sentences into SQL queries. Leveraging manually annotated datasets like SPIDER (Yu et al., 2019), numerous methodologies have emerged, including SpCQL (Guo et al., 2022) and SQLNet (Xu et al., 2017). Conversely, scant attention has been directed towards the Text2Cypher task. Existing approaches typically resort to decomposing a complete query into smaller components and translating each part separately. For instance, $R^{3}$-NL2GQL (Zhou et al., 2023b) partitions the query generation process into CRUD keywords prediction, clause selection, and object type identification. Despite the success of these methods, adapting them to a specific KG database demands substantial extra effort. With the rise of LLMs, using LLMs for Cypher query generation appears promising. Notably, to the best of our knowledge, no endeavors have explored the potential application of LLMs to the Text2Cypher task. Our work aims to bridge this gap in the literature.
@@
The Cypher writing performance of vanilla LLMs is not satisfactory. To improve it, we employ SFT, which necessitates a dataset of Question-Cypher pairs. However, creating such a dataset is challenging as it requires both domain-specific knowledge of the KG’s content and expertise in Cypher’s syntax. Consequently, there is currently no annotated dataset for the Text2Cypher task. To overcome this obstacle, we introduce SyntheT2C, a method designed to produce high-quality synthetic Question-Cypher pairs through two distinct pipelines: LLM-based prompting and template-filling (as shown in Figure 1). The LLM-based prompting pipeline aims to generate Cypher queries with greater semantic flexibility, while the template-filling pipeline focuses on producing syntactically complex Cypher queries. The generated Question-Cypher pairs undergo rigorous automated and manual validation, before being used to fine-tune backbone LLMs. The performance of Cypher generation is evaluated with a manually annotated evaluation dataset, complemented by a qualitative assessment using GPT as a judge. Additionally, we conduct a scalability test by fine-tuning the LLMs with larger synthetic datasets, which demonstrates that the synthetic data generated using our method does not collapse into simple patterns, thereby establishing the robustness of our approach for larger-scale applications.
@@
SyntheT2C is tested with two medical KG databases: the LHY database and the Hetionet database (details in Section 5). The generated synthetic dataset, “MedT2C”, will be made public.
@@ 2.1 Knowledge Graph and graph database
KGs have emerged as fundamental resources for organizing, representing, and querying vast amounts of interconnected information or domain-specific knowledge. These graphs find applications across various domains, including but not limited to, healthcare (Cui et al., 2024; Abu-Salih et al., 2022), finance (Elhammadi et al., 2020; Kertkeidkachorn et al., 2023), and e-commerce (Xu et al., 2020). In the realm of Natural Language Processing, KGs serve as invaluable sources of context and factual knowledge, enabling systems to reason, infer, and generate responses with enhanced accuracy and coherence. To handle the processing of graph data, a series of graph databases were invented, including Neo4j (Neo4j, 2012), NebulaGraph (Wu et al., 2022), and Amazon Neptune (Bebee et al., 2018). Among them, our work focuses on the Neo4j database (Neo4j, 2012), a widely adopted graph database management system. Neo4j database employs Cypher query language for expressing complex graph patterns during the retrieval.
@@ 2.2 Large Language Models
LLMs are advanced AI models that have been trained on vast amounts of text data to understand and generate human-like language. Following the milestone release of InstructGPT (Ouyang et al., 2022) by OpenAI, a series of LLMs are built, featuring different advantages and drawbacks, e.g., the series of GPT models (Brown et al., 2020; OpenAI, 2023) by OpenAI, Llama (Meta, 2024) by Meta, Qwen (Bai et al., 2023) by Alibaba Cloud, InternLM (Cai et al., 2024b) by Shanghai AI Lab, etc. Recent researches highlight LLMs’ ability to utilize external existing tools like calculator, search engine, or databases (Patil et al., 2023; Nakano et al., 2022; Cai et al., 2024a; Qin et al., 2023), which is usually abstracted as “Function calling”. Many of its implementations involve generating codes or queries to interact with external tools.
@@ 2.3 Code generation
Code Generation is the process of automatically producing executable code from a higher-level representation or natural language. With the advent of LLMs, code generation has experienced a significant advancement. LLMs can now be trained on vast amounts of code and programming-related text materials, enabling them to understand and generate code snippets based on given requirements (e.g., Codex (Chen et al., 2021), Polycoder (Xu et al., 2022), and Code Llama (Rozière et al., 2024)). By leveraging the contextual understanding (Dong et al., 2023) and language capabilities of LLMs, code generation becomes more efficient, accurate, and adaptable. Code generation with LLM is not only useful in helping developers to write codes but also in providing a powerful “language” for LLM to interact with other tools: LLMs can be tuned to output executable codes or queries to manipulate external resources. This is the fundamental idea for research in “Function Calling” and Multi-Agent Systems. Current code generation methods rely on two methods for evaluation: either with automatic metrics calculated with an annotated evaluation dataset (Papineni et al., 2002; Lin, 2004; Banerjee and Lavie, 2005; Evtikhiev et al., 2023; Zhou et al., 2023a) or with comparison by a judge (human or powerful LLM like GPT-4) (Zheng et al., 2023). Both evaluation methods are used in our work.
@@ 3.2.1 LLM-based prompting pipeline
This pipeline adopts an idea similar to Knowledge Distillation: we use the Cyphers generated by a stronger LLM to SFT weaker LLMs. Half of $\mathcal{S}$ is built by few-shot prompting GPT-4o (OpenAI, 2023). To simplify the process and ensure a higher quality of the generated data, we split the whole generation task into (1) extracting information from the database; (2) determining the question categories; and (3) generating the Cyphers for each category with extracted information.
@@ 3.2.2 Template-filling pipeline
One such template is illustrated in Figure 4. In this example, the subschema is introduced to manage cases where the entire database cannot be loaded at once, necessitating the selection and injection of only the relevant subgraph into the prompt. The variables label_i and prop_j represent the randomly sampled names of nodes and their attributes. These templates are initially crafted taking inspiration from Cypher Generator (Onofrei, 2024), then enriched and verified by the authors. More details about the construction process of the templates are presented in Appendix C.
@@ 3.3 Quality validation
To ensure the quality of the generated synthetic Question-Cypher pairs before their application in SFT, it’s imperative to conduct thorough validation. However, manually scrutinizing thousands of Cypher queries is arduous and time-consuming. In response, a suite of automatic validators are proposed to alleviate the burden of manual inspection. In the end, the Cyphers that pass through these automated validators undergo a final round of meticulous manual validation by researchers.
@@ 3.3.1 Automatic validation
The design of Semantic Validator is inspired by the research in machine translation (Hoang et al., 2018). This validator utilizes an LLM to translate the generated Cypher back into a natural language question. It then computes the semantic similarity between the translated question and the original question. If the similarity score exceeds a predefined threshold, the Cypher passes validation. We also tested an alternative implementation, where the LLM assesses semantic similarity directly. Both versions produce coherent validation results, with the latter being adopted for efficiency in subsequent experiments. The prompt used in this validator is presented in Appendix D.1.
The Entity Validator assesses the coverage of entities in the generated Cyphers. The entities in the original question $q$ are extracted via Named Entity Recognition (NER) using the spaCy (Honnibal and Montani, 2017) model en_core_web_sm. Entities in the generated Cypher $c$ are parsed and extracted using Regular Expressions. A successful validation requires 100% coverage of $q$’s entities in $c$. English entities are first transformed into lemmas using spaCy for fuzzy matching.
@@
Lastly, the Coherence Validator executes the Cypher against $\mathcal{B}$ and evaluates the coherence between the execution results and the original question with LLM (with prompt in Appendix D.2).
@@ 3.3.2 Manual validation
Each Cypher checked by the validators is randomly assigned to two researchers, who independently assess its quality. If both researchers provide a unanimous judgment, their consensus is adopted. In cases of divergent opinions, a third researcher is brought in for further review. The final validation outcome for such Cyphers is determined through a majority vote among the three researchers. Over 98% of the pairs passed the manual validation. Manual expertise involved is marginal as only less than 2% of the pairs failed the manual validation.
@@ 4.1 LHY and Hetionet Graph databases
In our experiment, we employed two Neo4j databases of general medical knowledge: the LHY Medical Knowledge Database (referred to as “LHY”) and the Hetionet Medical Knowledge Database (referred to as “Hetionet”). Both databases are publicly accessible, differing primarily in language: LHY is written in Chinese, whereas Hetionet is written in English. Their detailed statistics are presented in Appendix E.
The LHY Database (Liu, 2018) serves as the backend database for a Medical Question-Answering system. This database comprises comprehensive medical knowledge, encompassing a wide array of diseases, symptoms, drugs, and related information. Its content is sourced from medical websites, meticulously cleaned, reorganized, and stored within a Neo4j database. There are about 44k entities and 300k relations in it.
Hetionet (Himmelstein et al., 2017) is an open and free-to-use database of biomedical knowledge resource implementing “hetnet” model. Aggregating insights from 29 public databases, Hetionet boasts a knowledge network spanning various fields, encompassing a wide array of entities, including genes, compounds, anatomical structures, diseases, symptoms, side effects, etc. There are approximately 47k entities and 2.2 million relations in the Hetionet database.